Solve the following initial value problem

  1. Ut(x,t)=10Uxx(x,t) -1<x<1, t>0
  2. U(-1,t)=U(1,t) Ux(-1,t)=Ux(1,t) t>0
  3. Ux(x,0)=x+1 -1<x<1
1
Expert's answer
2021-12-27T04:19:16-0500

solution of heat equation:

"u(x,t)=\sum B_n sin(n\pi x/L)e^{-k(n\pi /L)^2t}"

where

"B_n=\frac{2}{L}\int^L_0f(x)sin(n\pi x/L)dx"

"f(x)=u(x,0)"

we have:

"k=10,L=1"

then:

"f(x)=\int^1_{-1}u_x(x,0)dx=\int^1_{-1}(x+1)dx=2"

"B_n=4\int^1_0sin(n\pi x)dx=-\frac{4}{\pi n}cos(n\pi x)|^1_0=-\frac{4}{\pi n}cos(n\pi )+\frac{4}{\pi n}"

"B_n=0" for even n

"B_n=\frac{8}{\pi n}" for odd n


"u_n(x,t)=0" for even n

"u_n(x,t)=\frac{8}{\pi n}sin(n\pi x)e^{-10(n\pi )^2t}" for odd n


for "u(-1,t)=u(1,t)" :

"u(-1,t)=\frac{8}{\pi n}sin(-n\pi )e^{-10(n\pi )^2t}=0"

"u(1,t)=\frac{8}{\pi n}sin(n\pi )e^{-10(n\pi )^2t}=0"


for "u_x(-1,t)=u_x(1,t)" :

"u_x(x,t)=8cos(n\pi x)e^{-10(n\pi )^2t}"

"u_x(-1,t)=8cos(-n\pi )e^{-10(n\pi )^2t}=-8e^{-10(n\pi )^2t}"

"u_x(1,t)=8cos(n\pi )e^{-10(n\pi )^2t}=-8e^{-10(n\pi )^2t}"